# Towards Domain Adaptive Neural Contextual Bandits

### v1

@@ 1 Introduction
Contextual bandit (CB) algorithms have shown great promise for naturally handling exploration/exploitation trade-off problems with optimal statistical properties. Notably, LinUCB [1] and its various adaptations [2, 3, 4, 5, 6, 7, 8, 9], have been shown to be able learn the optimal strategy when all data come from the same domain. However, these methods fall short when applied to data from a new domain. For example, a drug reaction prediction model trained by collecting feedback from mice (the source domain) may not work for humans (the target domain).
@@ 2 Related Work
Contextual Bandits. In the realm of adaptive decision-making, contextual bandit algorithms, epitomized by LinUCB [1], have carved a niche in efficiently balancing the exploitation-exploration paradigm across a spectrum of use cases, notably in complex adaptive systems such as recommender systems [1, 10]. These algorithmic frameworks, along with their myriad adaptations [2, 3, 4, 7, 5, 6, 9, 8, 11], have decisively outperformed conventional bandit models that operate devoid of contextual awareness [12]. This superiority is underpinned by theoretical models, paralleling the insights in [12], where LinUCB variants are validated to conform to optimal regret boundaries in targeted scenarios [13]. However, a discernible gap in these methodologies is their reduced efficacy in a new domain, particularly when getting feedback involves high cost. Our proposed DABand bridges this gap by adeptly aligning both source-domain and target-domain representations in the latent space, thereby reducing performance drop when transfer across different domains.
@@
Domain Adaptation. The landscape of domain adaptation has been extensively explored, as evidenced by a breadth of research [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. The primary objective of these studies has been the alignment of source and target domain distributions to facilitate the effective generalization of models trained on labeled source data to unlabeled target data. This alignment is conventionally attained either through the direct matching of distributional statistics [15, 26, 27, 20, 24] or via the integration of an adversarial loss [28, 29, 30, 19, 31, 21, 22, 23]. The latter, known as adversarial domain adaptation, has surged in prominence, bolstered by its theoretical foundation [32, 33, 19, 34], the adaptability of end-to-end training in neural network architectures, and its empirical efficacy. However, these methods only work in offline settings, and assumes complete observability of labels in the source domain. Therefore they are not applicable to our online bandit settings.
@@
Domain Adaptation Related to Bandits. There is also work related to both domain adaptation and bandits. Specifically, Guo et al. [35] propose a domain adaptation method using a bandit algorithm to select which domain to use during training.
@@
We note that their goal and setting is different from our DABand’s. (1) [35] focuses on improving accuracy in a typical, offline domain adaptation setting. In contrast, our DABand focuses on minimizing regret in an online bandit setting. (2) [35] assumes complete access to ground-truth labels in the source domain, while in our bandit setting, ground-truth may still be unknown if the action is incorrect. Therefore their work is not applicable to our setting.
@@ 3.1 Preliminaries
LinUCB. LinUCB [1] is a classic method for the typical contextual bandit setting. It assumes that the reward is linear to its input context, i.e., ${\bf x}_{i,a}$. During each round $i$, the agent selects an arm according to historical contexts and their corresponding rewards, $r_{i,a}$:
@@
where $A_{i-1}=\gamma{\bf I}_{d}+\sum\nolimits_{j=1}^{i-1}{\bf x}_{j,\widehat{a}_{j}}{\bf x}_{j,\widehat{a}_{j}}^{T}$ with $\gamma>0$ and ${\bf x}_{j,\widehat{a}_{j}}^{T}$ as the historical selected action’s context. ${\bf x}_{i,a}$ is the context for the candidate action $a$ in round $i$. $\alpha>0$ is a hyperparameter that adjusts the exploration rate.
LinUCB with Representation Learning. In this paper, we use Neural LinUCB [11, 10] as a backbone model to handle high-dimensional context with representation learning. We assume that there exist a ground-truth encoder $\phi^{*}$ to encode the raw context ${\bf x}_{i,a}$ into a latent-space representation (encoding) $\phi^{*}({\bf x}_{i,a})$. Subsequently, the reward for the context ${\bf x}_{i,a}$ is then $r({\bf x}_{i,a})=\langle\theta^{*},\phi^{*}({\bf x}_{i,a})\rangle$ plus some stochastic noise $\epsilon$. Furthermore, we use the same setting as in [11] where the ground-truth rewards are restricted to the range of $[0,1]$. Additionally, we further bound predicted rewards by normalizing $\|\widehat{\theta}\|=1$ and $\|\widehat{\phi}({\bf x}_{i,a})\|=1$.
In the single-domain typical setting, the goal is to learn an encoder $\widehat{\phi}$ and the contextual bandit parameter $\widehat{\theta}$, such that our estimated reward $\widehat{{r}}({\bf x}_{i,a})=\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a})\rangle$ can be close to the ground-truth reward $r({\bf x}_{i,a})$. One can then use this to form derive policies to minimize the regret in Eqn. 1.
@@ 3.2 Our Cross-Domain Contextual Bandit Setting
Difference between Eqn. 1 and Eqn. 2. In the typical setting, Eqn. 1 uses different updated policies for each round; this is also true for the source domain. In contrast, the target regret in Eqn. 2 use the same fixed policy obtained from the source domain for all $N$ (hypothetical) rounds because one cannot collect feedback from the target domain. See the difference between Definition 3.4 and Definition 3.5 below for a more formal comparison.
@@ 3.5 Cross-Domain Error Bound for Regression
Prior to introducing our final regret bound, it is necessary to define an additional component. A fundamental challenge in general domain adaptation problems is to manage the divergence between source and target domains. Unfortunately, previous domain adaptation theory only covers classification tasks [36, 19]. In contrast, the problem of contextual bandits is essentially a reward regression problem. To address this challenge, we introduce the following new definition to bound the error between the source and target domains.
@@
3:  Initialization: $A_{0}=\gamma{\bf I},b_{0}={\bf 0}$, and for $\widehat{\theta}_{0}$ and $\widehat{\phi}_{0}$ are initialized following [11].
@@
6:     Choose an action $\widehat{a}_{i}=\mathop{\rm argmax}_{a\in[K]}\left[\widehat{\phi}_{i-1}({\bf x}_{i,a}^{\mathcal{S}})\right]\widehat{\theta}_{i-1}+\alpha\|\left[\widehat{\phi}_{i-1}({\bf x}_{i,a}^{\mathcal{S}})\right]\|_{A_{i-1}}$.
@@
12:           Calculate $\mathcal{L}^{DABand}$ in Eqn. 10.
13:           Update encoder $\widehat{\phi}_{i}$ and discriminator $g$ by back-propagation to solve the minimax optimization in Eqn. 9.
@@
3.6 Final Regret Bound
Theorem 3.1 (Target Regret Bound).
$\leq{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum_{i=1}^{N}\Bigg{(}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}}_{Predicted~{}Reward}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}},$ (5)
where $\mathit{R}_{\mathcal{S}}$, $\epsilon_{\mathcal{S}}(h)$, and $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$ are defined in previous sections. $\psi$ and $C$ are constants which are independent of the problems.
  • (1)
  • (2)
  • (3)
    Typical generalization bounds in DA naively minimize the source regret $R_{{\mathcal{S}}}$ while aligning source and target data in the latent space (i.e., minimizing $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$). This is not sufficient, as we need two additional terms, i.e., the regression error $\epsilon_{\mathcal{S}}(h)$ and the predicted reward $\sum\nolimits_{i=1}^{N}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}$, as regularization terms.
  • (4)
  • (5)
    Minimizing the predicted reward $\sum\nolimits_{i=1}^{N}\Big{|}\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}$ regularizes the model to avoid overestimating rewards.
  • Regression Error in the Source Domain. $\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{S}}^{*},\phi_{\mathcal{S}}^{*}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}$ (i.e., $\epsilon_{\mathcal{S}}(h)$ in Theorem 3.1), which defines the difference between the true reward from selecting action $\widehat{a}_{i}$ and the estimated reward for this action.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. Details for more analysis are in Appendix F.3.1.
4 Method
With Theorem 3.1, we can then design our DABand algorithm to obtain optimal target regret. We discuss how to translate each term in Eqn. 5 to a differential loss term in Sec. 4.1 and then put them together in Sec. 4.2.
4.1 From Theory to Practice: Translating the Bound in Eqn. 5 to Differentiable Loss Terms
Source Regret. Inspired by Neural-LinUCB [11], we use LinUCB [1] to update $\widehat{\theta}$ and the following empirical loss function when updating the encoder parameters $\widehat{\phi}$ by back-propagation in round $i$:
Insights. From Eqn. 6, a fundamental difference between classification problems and bandit problems becomes apparent. In classification tasks, both the ground-truth label and the estimated label are accessible. However, in bandit problems, we only know the estimated label and reward. If the estimated rewards align with our estimated label’s correctness, it suggests knowledge of the ground-truth label for that feature. If not, we can only deduce that the estimated label is inaccurate, without identifying the correct label among the other candidates. This uncertainty significantly amplifies the complexity of contextual bandit problems.
@@ 4 Method
Data Divergence. The data divergence term in Eqn. 5 leads to following loss term:

### v2

@@ 1 Introduction
Contextual bandit (CB) algorithms have shown great promise for naturally handling exploration/exploitation trade-off problems with optimal statistical properties. Notably, LinUCB (Li et al., 2010) and its various adaptations (Yue & Guestrin, 2011, Agarwal et al., 2014, Li et al., 2016, Kveton et al., 2017, Foster et al., 2018, Korda et al., 2016, Mahadik et al., 2020, Zhou et al., 2019), have been shown to be able learn the optimal strategy when all data come from the same domain. However, these methods fall short when applied to data from a new domain. For example, a drug reaction prediction model trained by collecting feedback from mice (the source domain) may not work for humans (the target domain).
@@ 2 Related Work
Contextual Bandits. In the realm of adaptive decision-making, contextual bandit algorithms, epitomized by LinUCB (Li et al., 2010), have carved a niche in efficiently balancing the exploitation-exploration paradigm across a spectrum of use cases, notably in complex adaptive systems such as recommender systems (Li et al., 2010, Wang et al., 2022). These algorithmic frameworks, along with their myriad adaptations (Yue & Guestrin, 2011, Agarwal et al., 2014, Li et al., 2016, Korda et al., 2016, Kveton et al., 2017, Foster et al., 2018, Zhou et al., 2019, Mahadik et al., 2020, Xu et al., 2020), have decisively outperformed conventional bandit models that operate devoid of contextual awareness (Auer, 2002). This superiority is underpinned by theoretical models, paralleling the insights in (Auer, 2002), where LinUCB variants are validated to conform to optimal regret boundaries in targeted scenarios (Chu et al., 2011). However, a discernible gap in these methodologies is their reduced efficacy in a new domain, particularly when getting feedback involves high cost. Our proposed DABand bridges this gap by adeptly aligning both source-domain and target-domain representations in the latent space, thereby reducing performance drop when transfer across different domains.
@@
Domain Adaptation. The landscape of domain adaptation has been extensively explored, as evidenced by a breadth of research (Pan & Yang, 2009, Pan et al., 2010, Long et al., 2018, Saito et al., 2018, Sankaranarayanan et al., 2018, Zhang et al., 2019, Peng et al., 2019, Chen et al., 2019, Dai et al., 2019, Wang et al., 2020, Nguyen-Meidine et al., 2021, Xu et al., 2023). The primary objective of these studies has been the alignment of source and target domain distributions to facilitate the effective generalization of models trained on labeled source data to unlabeled target data. This alignment is conventionally attained either through the direct matching of distributional statistics (Pan et al., 2010, Tzeng et al., 2014, Sun & Saenko, 2016, Peng et al., 2019, Nguyen-Meidine et al., 2021) or via the integration of an adversarial loss (Ganin et al., 2016b, Zhao et al., 2017, Tzeng et al., 2017, Zhang et al., 2019, Kuroki et al., 2019, Chen et al., 2019, Dai et al., 2019, Wang et al., 2020). The latter, known as adversarial domain adaptation, has surged in prominence, bolstered by its theoretical foundation (Goodfellow et al., 2014, Zhao et al., 2018, Zhang et al., 2019, Zhao et al., 2019), the adaptability of end-to-end training in neural network architectures, and its empirical efficacy. However, these methods only work in offline settings, and assumes complete observability of labels in the source domain. Therefore they are not applicable to our online bandit settings.
@@
Domain Adaptation Related to Bandits. There is also work related to both domain adaptation and bandits. Specifically, Guo et al. (2020) propose a domain adaptation method using a bandit algorithm to select which domain to use during training.
@@
We note that their goal and setting is different from our DABand’s. (1) Guo et al. (2020) focuses on improving accuracy in a typical, offline domain adaptation setting. In contrast, our DABand focuses on minimizing regret in an online bandit setting. (2) Guo et al. (2020) assumes complete access to ground-truth labels in the source domain, while in our bandit setting, ground-truth may still be unknown if the action is incorrect. Therefore their work is not applicable to our setting.
@@ 3.1 Preliminaries
LinUCB. LinUCB Li et al. (2010) is a classic method for the typical contextual bandit setting. It assumes that the reward is linear to its input context, i.e., ${\bf x}_{i,a}$. During each round $i$, the agent selects an arm according to historical contexts and their corresponding rewards, $r_{i,a}$:
@@
where $A_{i-1}=\gamma{\bf I}_{d}+\sum\nolimits_{j=1}^{i-1}{\bf x}_{j,\widehat{a}_{j}}{\bf x}_{j,\widehat{a}_{j}}^{T}$ with $\gamma>0$ and ${\bf x}_{j,\widehat{a}_{j}}^{T}$ as the historical selected action’s context, ${\bf x}_{i,a}$ is the context for the candidate action $a$ in round $i$, $\theta_{i}$ is the bandit parameter in round $i$ and $\alpha>0$ is a hyperparameter that adjusts the exploration rate.
LinUCB with Representation Learning. In this paper, we use Neural LinUCB Xu et al. (2020), Wang et al. (2022) as a backbone model to handle high-dimensional context with representation learning. We assume that there exist a ground-truth encoder $\phi^{*}$ to encode the raw context ${\bf x}_{i,a}$ into a latent-space representation (encoding) $\phi^{*}({\bf x}_{i,a})$. Subsequently, the reward for the context ${\bf x}_{i,a}$ is then $r({\bf x}_{i,a})=\langle\theta^{*},\phi^{*}({\bf x}_{i,a})\rangle$ plus some stochastic noise $\epsilon$. Furthermore, we use the same setting as in Xu et al. (2020) where the ground-truth rewards are restricted to the range of $[0,1]$. Additionally, we further bound predicted rewards by normalizing $\|\widehat{\theta}\|=1$ and $\|\widehat{\phi}({\bf x}_{i,a})\|=1$.
In the single-domain typical setting, the goal is to learn an encoder $\widehat{\phi}$ and the contextual bandit parameter $\widehat{\theta}$, such that our estimated reward $\widehat{{r}}({\bf x}_{i,a})=\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,a})\rangle$ can be close to the ground-truth reward $r({\bf x}_{i,a})$. One can then use this to form derive policies to minimize the regret in Eqn. (1).
@@ 3.2 Our Cross-Domain Contextual Bandit Setting
Difference between Eqn. (1) and Eqn. (2). In the typical setting, Eqn. (1) uses different updated policies for each round; this is also true for the source domain. In contrast, the target regret in Eqn. (2) use the same fixed policy obtained from the source domain for all $N$ (hypothetical) rounds because one cannot collect feedback from the target domain. See the difference between Definition 3.4 and Definition 3.5 below for a more formal comparison.
@@ 3.5 Cross-Domain Error Bound for Regression
Prior to introducing our final regret bound, it is necessary to define an additional component. A fundamental challenge in general domain adaptation problems is to manage the divergence between source and target domains. Unfortunately, previous domain adaptation theory only covers classification tasks Ben-David et al. (2010), Zhang et al. (2019). In contrast, the problem of contextual bandits is essentially a reward regression problem. To address this challenge, we introduce the following new definition to bound the error between the source and target domains.
@@
3.6 Final Regret Bound
Theorem 3.1 (Target Regret Bound).
$\leq{\color[rgb]{0,0.88,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.88,0}\pgfsys@color@cmyk@stroke{0.91}{0}{0.88}{0.12}\pgfsys@color@cmyk@fill{0.91}{0}{0.88}{0.12}\underbrace{\mathit{R}_{\mathcal{S}}}_{Source~{}Regret}}+{\color[rgb]{1,0,0.5}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0.5}\pgfsys@color@cmyk@stroke{0}{1}{0.50}{0}\pgfsys@color@cmyk@fill{0}{1}{0.50}{0}\underbrace{2\cdot\epsilon_{\mathcal{S}}(h)}_{Regression~{}Error}}+{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\underbrace{N\cdot\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}(\mathcal{S},\mathcal{T})}_{Data~{}Divergence}}+{\color[rgb]{0.15,1,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.15,1,0.67}\pgfsys@color@cmyk@stroke{0.85}{0}{0.33}{0}\pgfsys@color@cmyk@fill{0.85}{0}{0.33}{0}\underbrace{\psi+C}_{Constant}}$
$\quad+{\color[rgb]{1,0.58,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.58,0}\pgfsys@color@cmyk@stroke{0}{0.42}{1}{0}\pgfsys@color@cmyk@fill{0}{0.42}{1}{0}\underbrace{\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)}_{Predicted~{}Rewards}},$ (5)
where $\mathit{R}_{\mathcal{S}}$, $\epsilon_{\mathcal{S}}(h)$, and $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$ are the source regret, source-domain error, and ${\mathcal{H}\Delta\mathcal{H}}$ divergence defined in Definition 3.4,  Definition 3.3, and Definition 3.6, respectively. $\psi$ is a constant independent to the problem and and $C$ is a constant which can be ignored (see the Appendix for more details) .
  • (1)
  • (2)
  • (3)
    Typical generalization bounds in DA naively minimize the source regret $R_{{\mathcal{S}}}$ while aligning source and target data in the latent space (i.e., minimizing $\widehat{d}_{\mathcal{H}\Delta\mathcal{H}}$). This is not sufficient, as we need two additional terms, i.e., the regression error $\epsilon_{\mathcal{S}}(h)$ and the predicted reward $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$, as regularization terms.
  • (4)
  • (5)
    Minimizing the predicted reward $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$ regularizes the model to avoid overestimating rewards.
  • Regression Error in the Source Domain. $\sum\nolimits_{i=1}^{N}\Bigg{(}\Big{|}\langle\theta_{\mathcal{S}}^{*},\phi_{\mathcal{S}}^{*}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle-\langle\widehat{\theta},\widehat{\phi}(x_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\Big{|}\Bigg{)}$ (i.e., $\epsilon_{\mathcal{S}}(h)$ in Theorem 3.1) defines the difference between the true reward from selecting action $\widehat{a}_{i}$ and the estimated reward for this action.
  • Predicted Reward. $\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{T}})\rangle\right|\right)+\mathbbm{1}{[a_{i}^{*}\neq\widehat{a}_{i}]}\left(\sum^{N}_{i=1}\left(\left|\langle\widehat{\theta},\widehat{\phi}({\bf x}_{i,\widehat{a}_{i}}^{\mathcal{S}})\rangle\right|\right)\right)$ serves as a regularization term to regularize the model to avoid overestimating rewards.
The results of the ablation study in Table 4 in Sec. 5 highlight the significance of these two terms. Details for more analysis are in Appendix G.4.1.
4 Method
With Theorem 3.1, we can then design our DABand algorithm to obtain optimal target regret. We discuss how to translate each term in Eqn. (5) to a differential loss term in Sec. 4.1 and then put them together in Sec. 4.2.
4.1 From Theory to Practice: Translating the Bound in Eqn. (5) to Differentiable Loss Terms
Source Regret. Inspired by Neural-LinUCB (Xu et al., 2020), we use LinUCB Li et al. (2010) to update $\widehat{\theta}$ and the following empirical loss function when updating the encoder parameters $\widehat{\phi}$ by back-propagation in round $i$:
Insights. From Eqn. (6), a fundamental difference between classification problems and bandit problems becomes apparent. In classification tasks, both the ground-truth label and the estimated label are accessible. However, in bandit problems, we only know the estimated label and reward. If the estimated rewards align with our estimated label’s correctness, it suggests knowledge of the ground-truth label for that feature. If not, we can only deduce that the estimated label is inaccurate, without identifying the correct label among the other candidates. This uncertainty significantly amplifies the complexity of contextual bandit problems.
@@ 4 Method
3:  Initialization: $A_{0}=\gamma{\bf I},b_{0}={\bf 0}$, and for $\widehat{\theta}_{0}$ and $\widehat{\phi}_{0}$ are initialized following (Xu et al., 2020).
@@
6:     Choose an action $\widehat{a}_{i}=\mathop{\rm argmax}_{a\in[K]}\left[\widehat{\phi}_{i-1}({\bf x}_{i,a}^{\mathcal{S}})\right]\widehat{\theta}_{i-1}+\alpha\left\|\left[\widehat{\phi}_{i-1}({\bf x}_{i,a}^{\mathcal{S}})\right]\right\|_{A_{i-1}^{-1}}$.
@@
12:           Calculate $\mathcal{L}^{DABand}$ in Eqn. (10).
13:           Update encoder $\widehat{\phi}_{i}$ and discriminator $g$ by back-propagation to solve the minimax optimization in Eqn. (9).
@@
Data Divergence. The data divergence term in Eqn. (5) leads to following loss term: